OpenStudy (anonymous):

If an object is dropped from a height of 144 ft, the function h(t) = –16t² + 144 gives the height of the object after t seconds. When will the object hit the ground?

OpenStudy (anonymous):

ok

OpenStudy (mysesshou):

is this the whole thing? \[h(t) = -16 t^2 + 144\] because then t=0, which can't happen if you're dropping it 144ft a little confused. :(

OpenStudy (anonymous):

yes that is the height of the oblect AFTER t seconds. and the question asks when will the object hit the ground.

OpenStudy (mysesshou):

Ah, so perhaps the 144ft is not needed. The ground is when h=0, so that works to solve for \[0 = -16 t^2 + 144\] and solve for t ??

OpenStudy (mysesshou):

Do you think that works ?

OpenStudy (anonymous):

no find out how many seconds it will take for the object to hit the gorund

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

nvm i found the answer, its 3 seconds

OpenStudy (mysesshou):

sorry I had misunderstood. So, if it wants when it hits the ground, then I think h=0, giving: \[144 = -16 t^2 \] \[\sqrt{-144/(-16)} = t = 3 sec\] Do you think this works?

OpenStudy (mysesshou):

Sorry I wasn't much good help :(

OpenStudy (mysesshou):

good for you for finding it :)

OpenStudy (mysesshou):

sorry !